i am thinking of three positive numbers. the second number is five more than the first number, the third number is 3 less than 2 times the first number, the sum of all the numbers is 254, what are the three numbers?
Found 2 solutions by ptfile, TimothyLamb:
Answer by ptfile(81)

Let x = the first number
x+5 = the second number
2x-3 = the third number
The sum of all the numbers is 254.
x+x+5+2x-3=254
4x+2=254
4x=252
x=63
Substitute x for 63
63 + 5 = 68
2(63)-3= 123
The numbers are 63, 68 and 123.
Check
63 + 68 + 123 = 254
254 = 254

Answer by TimothyLamb(4379)

y = x + 5
z = 2x - 3
x + y + z = 254
---
put the system of linear equations into standard form
---
x - y = -5
2x - z = 3
x + y + z = 254
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 63
y = 68
z = 123
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---